OpenStudy (anonymous):

What is the sum of the first five terms of the geometric sequence in which a 1 = 8 and r = 1/4? Express your answer as an improper fraction using the slash (/) key and no spaces. Reduce the fraction to its lowest terms.

OpenStudy (anonymous):

\[8,8\times \frac{1}{4},8\times \frac{1}{4^2}, 8\times \frac{1}{4^3}\]

OpenStudy (anonymous):

8+1/4^x so adding on

OpenStudy (anonymous):

Sorry I got image fixed and I see it now,okay